through 3,4 m=-7/9

Answer by mananth(16946)

m= - 7/ 9

Plug value of the slope and point ( 3 , 4 ) in
Y = m x + b
4.00 = -7/3 + b
b= 4 - -7/3
b= 19/ 3
So the equation will be
Y = - 7/9 x + 19/3
m.ananth@hotmail.ca